# KeySpace: Public Key Infrastructure Considerations in Interplanetary Networks

### v1

@@ VII-C Key Revocation
For this scenario, we focus on the Earth/Moon, Earth/Mars, and Earth/Moon/Mars networks, with Iridium as a reference terrestrial satellite network. We ran 1 04410441\,0441 044 simulations with different configurations, discussing the most interesting results in this paper. [^9] The full results are provided at deep-space-network-simulator.github.io.
Figure 7 illustrates the race between an attacker’s message and a revocation – when the two are located in different segments, the attacker’s segment takes a long time to be covered due to the long distances involved, but other segments may be protected significantly earlier. We can see this reflected in the results in Figure 8.101010Note that in these results, a negative coverage time means the node will be protected from messages created before the revocation, since the revocation will reach the CA before the attacker’s message. We see that when the attacker and revocation are based in the same segment, the time taken to cover nodes largely depends upon each node’s proximity to a CA, with fast overall revocation coverage but some nodes being covered later than others. Our goal is therefore to optimize the PKI configuration to provide coverage in each case as early as possible, minimizing damage within the speed-of-light constraints inherent to long-distance communication.
@@
We can see aggregated coverage information for a wider range of configurations in Tables IV (Iridium) and V (Earth/Moon/Mars). The coverage time in these tables is given as the time after which all nodes in the segment are protected by the revocation. We can immediately see a significant difference between the cases where the attack originates from the same segment as the revocation, and when it originates from a different segment – in the latter case it takes significantly longer for the segments to be covered. This is because the revocation does not always have time to propagate to the segment in which the attacker is present, and must instead race the attacker’s messages to the victim/CA.
@@
Similarly, we focus on the best-case scenario for the attacker, in which the nodes already have the attacker’s certificate status cached, so nodes will not immediately request up-to-date revocation information from the CA. The results for when caching is not used can be found in the full dataset.
@@
VII-C1 Subscribe
Having noted that a large number of messages get accepted in cases where the attack and revocation originate from different segments of the network, we evaluate some potential methods of mitigating this threat. One such method requires OCSP CAs to register subscriptions for nodes that request a certificate status, forwarding any updates or revocations directly to that node. [^11] Note that this is not possible for OCSP Stapling, as the CA does not know the destination of the stapled certificate status. This is good practice in general, as it can cause caches of compromised certificates to be removed from the victim node sooner, reducing the window of opportunity for the attacker.
@@ VII-C1 Subscribe
VII-C2 Firewall
Finally, we propose a new technique in which the relay nodes are used as a firewall to filter out messages which use revoked certificates. Under this configuration, the relay filters out all messages passing through which are making use of revoked keys, stopping them from ever reaching their destination. This does not affect the message’s path, since all interplanetary traffic travels via the relays anyway, and in a number of cases (OCSP with Validators, OCSP Hybrid) the certificates are also validated by these relays. The technique has the potential to reduce the race between the revocation and the attacker’s messages when messages are waiting to be forwarded at a relay, allowing all of these messages to be dropped by the firewall instead. It also allows relays to drop messages from an attacker even if a valid certificate status has already been stapled to the message, reducing the reach of the attacker across long-distance relays. [^12] Depending on implementation, the firewall may inform the recipient about dropped messages, or simply label the message as unauthenticated so the recipient can choose what to do with it.
The results of this can be seen in Table VII. We can see that this technique has a huge impact on the coverage time of the revocations, particularly when the attack and revocation originate from different segments – the attacker’s segment is largely unaffected, but the other segments in the network can be protected substantially earlier. This brings us closer to the expected best-case performance of revocations under distributed PKI, bounding the attacker’s reach to the smallest segment possible and covering the other segments in the network as early as possible. This technique also reduces load on the interplanetary relays, preventing denial of service by attackers using compromised keys to use up all the bandwidth on relay links.
@@ VII-C2 Firewall
VIII Discussion
We already know that there is no “one size fits all” approach to key management in interplanetary networks, and our results confirm this – distributed OCSP with firewalls on each of the relays can provide high performance with fast revocation coverage, but some higher security applications may require centralized control, even at the cost of significantly increased latency. Our results enable satellite operators to make informed decisions about PKI implementation based on their specific needs and constraints: DSNS can be used to test the specifics of a PKI system against a network topology before launch and deployment, ensuring the desired balance between security and performance can be achieved. Additionally, our testing framework can be used in the future (alongside DSNS or another network simulator) to test newly proposed key management systems against a wide range of network topologies, establishing its performance and security characteristics.
Future work may consider testing “optimistic” protocols – that is, protocols which initially assume the certificate to be valid, rolling back the state if this is later shown not to be the case – or protocols like QUIC, which enable communication with zero round-trip-time setup [iyengarQUIC2021]. However, this raises additional security concerns: firstly, the system must be designed in such a way that state can be efficiently rolled back without losing other data or leaving artifacts behind. Many satellites have tight resource constraints, and it is vital that adversaries are prevented from exploiting these resources without proper authorization by opening temporary communication channels with invalid credentials. If these protocols are to be used, there must still be a mechanism for establishing contracts such that resources are not wasted on unauthorized users. Similarly, future work may wish to evaluate the scope and limitations for retroactive revocations, considering which actions can be safely rolled back without affecting general operation.
@@ VIII Discussion
We must also consider the complexity of implementation of these technologies. Some works evaluated in Section III introduce new protocols and PKI systems; although this enables more significant changes, these new protocols have not been tested as extensively. By using protocols that have already been standardized and are in wide use across the terrestrial internet, we minimize the risk of security threats in the PKI system itself. We also benefit from the fact that standards already exist, so standardization bodies for space systems such as CCSDS do not need to build new standards from scratch, instead building off what already exists. This could involve further additions to the “BPSec” proposed standard [birraneBundle2022], adding new block definitions and behaviors to enable the new functionality. This offers the opportunity to accelerate the timeline for standardization, which is critical when the space industry is moving so quickly – a lack of standards for interplanetary communication when new systems are deployed could result in incompatibility between systems, making the goal of an interplanetary internet more difficult to realize.
IX Conclusion
In this paper we have designed a framework for testing key management systems in satellite networks – this enables protocol designers to ensure their system is efficient and secure, and allows operators to find the optimal configuration for their network and specific needs. We have tested this framework against terrestrial PKI techniques applied to interplanetary networks using a novel network simulator. In contrast to current consensus, we have shown that terrestrial PKI can be used to secure communication in current and future satellite networks, ensuring efficient, low-latency connection establishment by distributing certificate authorities across the network.
@@ Appendix A Experiment Configuration
In the simulations built in Section VI, the terrestrial portion of the network contains 256 ground stations. These are randomly placed on Earth’s land surface, with the positions remaining fixed between simulations. The positions of the ground stations are visualized in Figure 9.

### v3

@@ 7.3 Key Revocation
For this scenario, we focus on the interplanetary Earth/Moon/Mars network, with Iridium as a reference terrestrial satellite network. We ran 1 04410441\,0441 044 simulations with different configurations, discussing the most interesting results in this paper. [^8] The full results are provided at deep-space-network-simulator.github.io.
Figure 6 illustrates the coverage time for a revocation under OCSP. We can see that the coverage time is often negative for segments in which the attacker is not present – this is due to the long distances between segments, enabling the revocation to reach the CA long before the attacker’s message traverses the relay link. As a result, only messages sent significantly before of the revocation will be accepted in that segment, resulting in a negative coverage time. This is explained further in Appendix D.
In contrast, when the attacker and revoker are located in the same segment, coverage time is primarily influenced by each node’s proximity to the CA, leading to rapid overall coverage with some nodes protected later than others. Our objective is to optimize the PKI configuration to achieve the earliest possible coverage in each scenario, minimizing damage within the speed-of-light constraints inherent in long-distance communication.
@@
We can see aggregated coverage information for a wider range of configurations in Tables IV (Iridium) and V (Earth/Moon/Mars). In these tables, the coverage time is the time after which every node in the segment is protected by the revocation. We can once again see a significant difference between the cases where the attack originates from the same segment as the revocation, and when it originates from a different segment – in the latter case it takes significantly longer for the segments to be covered. This is because the revocation does not always have time to propagate to the segment in which the attacker is present, and must instead race the attacker’s messages to the victim/CA.
@@
Similarly, we focus on the best-case scenario for the attacker, in which nodes already have the attacker’s certificate status cached, so they will not immediately request up-to-date revocation information from the CA. The results for when caching is not used can be found in the full dataset.
@@
7.3.1 Subscribe
Having noted that a large number of messages get accepted in cases where the attack and revocation originate from different segments of the network, we evaluate some potential methods of mitigating this threat. One such method requires OCSP CAs to register subscriptions for nodes that request a certificate status, forwarding any updates or revocations directly to that node. This is not possible for certain configurations (e.g. OCSP Stapling, as the CA does not know the destination of its stapled certificate status) but is good practice in supported cases, as caches of compromised certificates can be cleared sooner, reducing the window of opportunity for the attacker.
@@ 7.3.1 Subscribe
7.3.2 Firewall
Finally, we propose a new technique in which the relay nodes are used as a firewall to filter out messages which use revoked certificates. Under this configuration, the relay filters out all messages passing through which are making use of revoked keys, stopping them from ever reaching their destination or marking them as invalid (depending on implementation). This does not affect the message’s path, since all interplanetary traffic travels via the relays anyway, and in a number of cases (OCSP with Validators, OCSP Hybrid) the certificates are also validated by these relays. The technique has the potential to reduce the race between the revocation and the attacker’s messages when messages are waiting to be forwarded at a relay, allowing all of these messages to be dropped by the firewall instead. It also allows relays to drop messages from an attacker even if a valid certificate status has already been stapled to the message, reducing the reach of the attacker across long-distance relays.
The results of this can be seen in Table VII. We can see that this technique has a huge impact on the coverage time of the revocations, particularly when the attack and revocation originate from different segments – the attacker’s segment is largely unaffected, but the other segments in the network can be protected substantially earlier. This brings us closer to the expected best-case performance of revocations under distributed PKI, bounding the attacker’s reach to the smallest segment possible and covering the other segments in the network as early as possible. This technique also reduces load on the interplanetary relays, preventing denial of service by attackers using compromised keys to exhaust the bandwidth of relay links.
8 Discussion
Related works hypothesize that there is no “one size fits all” approach to PKI in satellite networks [viswanathanArchitecture2016, templinDTN2016]. Our results further confirm this: distributed OCSP with firewalls on each of the relays can provide high performance with fast revocation coverage, but some higher security applications may require centralized control, even at the cost of significantly increased latency. Our results enable satellite operators to make informed decisions about PKI implementation based on their specific needs and constraints: DSNS can be used to test the specifics of a PKI system against a network topology before launch and deployment, ensuring the desired balance between security and performance can be achieved. Additionally, our testing framework can be used in the future (alongside DSNS or another network simulator) to test newly proposed key management systems against a wide range of network topologies, establishing its performance and security characteristics.
@@ 8 Discussion
We must also consider the complexity of implementation of these technologies. Some works evaluated in Section 3 introduce new protocols; although this enables more significant changes, these new protocols have not been tested as extensively. By using protocols that have already been standardized and are in wide use across the terrestrial internet, we minimize the risk of security threats in the PKI system itself. We also benefit from the fact that standards already exist, so standardization bodies such as CCSDS do not need to build new standards from scratch, instead building off what already exists. This could involve further additions to the “BPSec” proposed standard [birraneBundle2022], adding new block definitions and behaviors to enable the new functionality. This offers the opportunity to accelerate the timeline for standardization, which is critical when the space industry is moving so quickly – a lack of standards for interplanetary communication when new systems are deployed could result in incompatibility between systems, making the goal of an interplanetary internet more difficult to realize.
8.1 Future Work
Alongside standardization efforts, some open questions remain which could be addressed in future work. To test cryptosystems and other implementation-specific details, DSNS could be extended to support bandwidth and storage simulation and message loss. This will enable regulators to figure out the specifics of how PKI protocols should be implemented to provide the fast connection establishment and effective revocations demonstrated in this paper, while also remaining within the computational and storage limits of satellite hardware. It will also enable demonstration of the concrete benefits of federated constellations, making use of a small number of powerful satellites to provide connectivity to a large number of devices while using LEO satellites for short-distance connections with low latency.
Future research may also consider testing “optimistic” protocols – that is, protocols which initially assume the certificate to be valid, rolling back the state if this is later shown not to be the case – or protocols like QUIC, which enable communication with zero round-trip-time setup [iyengarQUIC2021]. However, this raises additional security concerns: firstly, the system must be designed in such a way that state can be efficiently rolled back without losing other data or leaving artifacts behind. Many satellites have tight resource constraints, and it is vital that adversaries are prevented from exploiting these resources without proper authorization by opening temporary communication channels with invalid credentials. If these protocols are to be used, there must still be a mechanism for establishing contracts such that resources are not wasted on unauthorized users. Similarly, there is scope for evaluation of the scope and limitations for retroactive revocations, considering which actions can be safely rolled back without affecting general operation.
9 Conclusion
In this paper we have designed a framework for testing key management systems in satellite networks – this enables protocol designers to ensure their system is efficient and secure, and allows operators to optimize configurations for their network and specific needs. We have tested this framework against terrestrial PKI techniques applied to interplanetary networks using a novel network simulator. In contrast to current consensus, we have shown that terrestrial PKI can be used effectively in current and future satellite networks, ensuring efficient, low-latency connection establishment by distributing certificate authorities across the network.
@@ Appendix A Experiment Configuration
In the simulations built in Section 6, the terrestrial portion of the network contains 256 ground stations. These are randomly placed on Earth’s land surface, with the positions remaining fixed between simulations. The positions of the ground stations are visualized in Figure 7; exact coordinates will be included in the experiment configuration code released on publication (see Appendix E).
Appendix B Protocols
In this appendix we give sequence diagrams for each of the PKI techniques tested in this paper. These diagrams are shown for the distributed case, in which there are multiple CAs and authority is delegated. Note that the diagrams are also for nominal operation only – they do not illustrate behavior in the case of invalid signatures, caching, or revocations. OCSP Hybrid is not shown here, since behavior is identical to OCSP Stapling for communication within a segment, and identical to OCSP with Validators for communication between segments.
Full protocol specifications are in the source code for DSNS.
OCSP
ABCABSignedMessageCertificateQueryMessageCertificateStatusMessage
OCSP Stapling
ABCAACertificateQueryMessageCertificateStatusMessage StapledMessage ( SignedMessage, CertificateStatusMessage )
OCSP with Validators
ACARelayB ValidatorQueryMessage ( SignedMessage ) StapledMessage ( SignedMessage, CertificateStatusMessage )
CRL
ABCAASignedMessageCRLRequestMessageCRLDirectMessage
CRL Broadcast
ACAACRLBroadcastMessageCRLBroadcastMessageCRLBroadcastMessage
Appendix C Latency Results
In this appendix we include additional results on the point-to-point latency of terrestrial constellations. Figure 9 illustrates how the latency changes over time for the Iridium and Starlink constellations, and the federated constellation of CubeSats evaluated in the paper.
Appendix D Revocation Coverage
In this appendix we provide further explanation of the “coverage time” results given for the revocation experiments in Section 7.3.
Figure 8 illustrates the race between an attacker’s message and a revocation, when they originate from different segments. Due to the vast distances involved, each segment is covered rapidly by its respective message, but cross-segment messages experience a significant delay. Consequently, the attacker must send their message well in advance of the revocation to ensure it is accepted, as the revocation message will reach the victim CA much more quickly than the attacker’s message. This can result in a negative coverage time, where messages sent before the revocation are rejected in the segment from which the revocation originates, while messages sent after the revocation are still accepted in the attacker’s segment.
Appendix E Data Availability
In the course of writing this paper, we ran a large number of network simulations, each of which produced substantial logs. These will be made fully available on publication via institutional research archives. Metrics for each simulation following processing can be found at deep-space-network-simulator.github.io.
Additionally, the source code for DSNS will be made available on publication as a Git repository, and released under a permissive license to enable future research. This will include the code used to configure and run the experiments, and process the results.